OpenStudy (anonymous):

Which answer shows this equation written in standard form? 4 – 3(x – y) = –8x + 2

OpenStudy (anonymous):

A. –11x + 3y = –2 B. 5x – 3y = 2 C. 5x + 3y = –2 D. 5x + y = 2

OpenStudy (anonymous):

What do you think the answer is?

OpenStudy (anonymous):

b

OpenStudy (anonymous):

Um i would say C

OpenStudy (anonymous):

oh ok